# JWST captures a sudden stellar outburst and inner disk wall destruction

### v1

@@ 5 Summary
In this study, we have acquired and analyzed the JWST/MIRI-MRS spectrum of T Cha, an accreting G8 star surrounded by a transition disk, exhibiting highly variable optical emission. This disk is known for its large dust gap separating the small inner disk ($<$1 au) and the outer disk ($\sim$ 20-50 au). The MIRI-MRS spectrum of T Cha shows an unexpected change (seesaw variability) compared to a previous Spitzer spectrum and all other photometric points observed before 2020. The emission at shorter wavelength ($\lambda<10\mu m$) decreases while at longer wavelength ($\lambda>10\mu m$) increases, both by a factor of up to 3. This mid-IR change is consistent with recent WISE observations and accompanied with the less variable but brighter optical emission during 2021-2023 period. We use the radiative transfer code RADMC-3D to model both the Spitzer and JWST spectra, and find that both the mid-IR and optical behavior of T Cha can be explained by the optically thick asymmetric inner disk wall significantly decreased in mass and height, and possibly due to an outburst.
@@
Our observations and models suggest that T Cha is experiencing episodic event (possible accretional outburst) every 5-10 years. In each event, an appreciable amount of micron-sized dust grains are removed from the inner disk. This suggests that T Cha will show a significant change in its SED in the following decades due to the decrease and depletion of the micron-sized dust grains, and will be an excellent system to study inner disk dust replenishment and possibly the formation of second generation dust. We propose that future high spatial and spectral resolution observations including spectra and images at different bands will help us understand the last stage of evolution of transition disks.
@@
This work is based on observations made with the NASA/ ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. Some/all of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi:10.17909/dhmh-fx64 (catalog DOI). The observations are associated with JWST GO Cycle 1 program ID 2260. This work was funded by NASA/STScI GO grant JWST-GO-02260.001. We thank Andras Gaspar and Jane Morrison for their efforts in acquiring and processing the data. This work was also supported by the NKFIH excellence grant TKP2021-NKTA-64. C.X. and I.P. acknowledge partial support by NASA under Agreement No. 80NSSC21K0593 for the program “Alien Earths”. A.D.S. was supported by the European Union’s Horizon 2020 research and innovation program under Marie Sklodowska-Curie grant agreement No. 823823 (DUSTBUSTERS). G.B. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agree- ment No. 853022, PEVAP). A.D.S. acknowledges the support of a Science and Technology Facilities Council (STFC) PhD studentship and funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 1010197S1 MOLDISK). R.A. acknowledges funding from the Science & Technology Facilities Council (STFC) through Consolidated Grant ST/W000857/1. C.X. thanks Chia-Lung Lin for the useful discussion on the variability mechanisms of variable stars. We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work is based in part on archival observations made with the Spitzer Space Telescope, which was operated by the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA.
@@ Appendix B Photometry Table
Table 5: References of the photometric points

### v3

@@ 5 Summary
In this study, we have acquired and analyzed the JWST/MIRI-MRS spectrum of T Cha, an accreting G8 star surrounded by a transition disk, exhibiting highly variable optical emission. This disk is known for its large dust gap separating the small inner disk ($<$1 au) and the outer disk ($\sim$ 20-50 au). The MIRI-MRS spectrum of T Cha shows an unexpected change (seesaw variability) compared to a previous Spitzer spectrum and all other photometric points observed before 2020. The emission at shorter wavelength ($\lambda<10\mu m$) decreases while at longer wavelength ($\lambda>10\mu m$) increases, both by a factor of up to 3. This mid-IR change is consistent with recent WISE observations and accompanied with a less variable but brighter optical emission during 2021-2023. We use the radiative transfer code RADMC-3D to model the Spitzer and JWST spectra along with the entire SED, and find that both the mid-IR and optical behavior of T Cha can be explained by an optically thick asymmetric inner disk wall with significantly decreased mass and height, possibly due to an outburst.
@@
Our observations and models suggest that T Cha is experiencing episodic events, possibly accretion outbursts, every 5-10 years. In each event, an appreciable amount of micron-sized dust grains are removed from the inner disk. This suggests that T Cha may show a significant change in its SED in the following decades due to the decrease and depletion of the micron-sized dust grains, and will be an excellent system to study inner disk dust replenishment. Future high spatial and spectral resolution observations at different bands can help understand the last stage of evolution of this and other transition disks.
@@
This work is based on observations made with the NASA/ ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. Some/all of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/dhmh-fx64 (catalog doi:10.17909/dhmh-fx64). The observations are associated with JWST GO Cycle 1 program ID 2260. This work was funded by NASA/STScI GO grant JWST-GO-02260.001. We thank Andras Gaspar and Jane Morrison for their efforts in acquiring and processing the data. This work was also supported by the NKFIH excellence grant TKP2021-NKTA-64. C.X. and I.P. acknowledge partial support by NASA under Agreement No. 80NSSC21K0593 for the program “Alien Earths”. A.D.S. was supported by the European Union’s Horizon 2020 research and innovation program under Marie Sklodowska-Curie grant agreement No. 823823 (DUSTBUSTERS). G.B. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agree- ment No. 853022, PEVAP). A.D.S. acknowledges the support of a Science and Technology Facilities Council (STFC) PhD studentship and funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 1010197S1 MOLDISK). R.A. acknowledges funding from the Science & Technology Facilities Council (STFC) through Consolidated Grant ST/W000857/1. C.X. thanks Chia-Lung Lin for the useful discussion on the variability mechanisms of variable stars. We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work is based in part on archival observations made with the Spitzer Space Telescope, which was operated by the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA.
@@ Appendix A Connection between parameter changes and SED changes
In addition, we also present here the influence on the SED from adding small Silicate grains in the inner disk (see Fig 8). In this model, we remove the inner disk wall, and leave the previously hidden mass with the same composition as the outer disk (70% of silicate and 30% of graphite) as the main emission source of the NIR wavelength. We vary other parameters to get a relatively good fit with our JWST spectrum. However, the strong 10 $\mu$m Si feature in this case does not match our data, meaning that the inner disk wall cannot be completely destroyed.
Figure 8: Model for inner disk without a wall (purple) compared with the model of a wall $+$ hidden mass. We can see that the model without a wall will have a strong 10 $\mu$m silicate feature, which does not match our observed data.
@@ Appendix B Photometry Table
Table 5: References of the photometric points. The photometric points used to model the SED (see Fig. 3) can also be downloaded here.